# Spectral Clustering with Likelihood Refinement is Optimal for Latent Class Recovery

### v1

@@ 4 Theoretical Guarantees
$\theta_{j,k}\in[c_{\theta},C_{\theta}]$ for some universal constants $10<c_{\theta}<C_{\theta}<$ 1, and $j\in[J]$, $k\in[K]$.
@@ 4.1 Theoretical Guarantee of Spectral Clustering
Proposition 1 (Adapted from Theorem 3.1 in Zhang and Zhou, (2024)).
Define $\sigma^{2}_{\theta}:=\left(1-2\theta\right)/\left[2\log\left(\left(1-\theta\right)\theta^{-1}\right)\right]$ for $\theta\in(0,1)\backslash\left\{1/2\right\}$ and $\sigma^{2}_{\theta}:=1/4$ for $\theta=1/2$, and let $\bar{\sigma}:=\max_{j,k}\sigma_{\theta_{j,k}}$. Assume $\text{rank}\left(\mathbf{\Theta}\right)=K$, $\alpha N/K^{2}\geq 10$ and
@@
Proposition 1 delivers that the simple and computationally efficient spectral clustering can essentially lead to exponentially small error rate with respect to the latent class separation $\Delta$. Notably, we can have exact recovery of all latent class labels, (that is, $\ell\left(\widetilde{\bm{s}},{\bm{s}}^{\star}\right)$ with $\widetilde{s}_{i}=\pi(s_{i})$ for all $i=1,\ldots,N$) with high probability as long as $\Delta^{2}/\bar{\sigma}^{2}\geq 8\left(1+c\right)\log N$. In the LCM we study, we have $\bar{\sigma}\leq 1/4$ by definition and hence it suffices to have $\Delta^{2}\geq 2(1+c)\log N$ to achieve exact recovery. This is not hard to achieve even for slowly growing number of items, because if we have constant separation of item parameters for each item with $\delta=\min_{k_{1}\neq k_{2}\in[K],j\in[J]}|\theta_{j,k_{1}}-\theta_{j,k_{2}}|$, then $\Delta^{2}=J\delta^{2}$. In this case, having $J>2(1+c)\delta^{-2}\log N$ for exact recovery corresponds to a moderately growing dimension. In sharp contrast, the JML estimator analyzed in Zeng et al., (2023) only achieves a polynomial convergence rate of order $O\left(J^{-1/2}\right)$. Moreover, they can not achieve exact recovery according to the technical condition and conclusion therein. Therefore, the new spectral estimator $\widetilde{\bm{s}}$ represents a substantial improvement with exponential error decay in (13).
@@
Despite the above nice theoretical guarantee for spectral clustering, the clustering error’s reliance on the Euclidean metric $\Delta$ in Proposition 1 is inherent from the general framework analyzed in Zhang and Zhou, (2024), and it turns out to be not statistically optimal due to ignorance of the likelihood information. We will discuss and address this issue next.
@@ 4.2 Fundamental Statistical Limit for LCMs
In this subsection, we provide the understanding of the information-theoretic limit (fundamental statistical limit) of clustering in LCMs. To compare different statistical procedures, we consider the minimax framework that had been widely considered in both statistical decision theory and information theory in the past decades (Le Cam,, 2012; Takezawa,, 2005). The minimax risk is defined as the infimum, over all possible estimators, of the maximum loss (here, the mis-clustering error) taken over the entire parameter space. In other words, it is the rate of the best estimator one can have among all statistical procedures in the worst-case scenario.
Formally, we establish the minimax lower bound for mis-clustering error under the fixed-effect LCM in the following result, whose proof is adapted from Lyu et al., (2025) and can be found in Section B.1.
@@
Roughly speaking, Theorem 1 informs us that under the fixed-effect LCM, the error rate $\exp\left(-I^{\star}/2\right)$ is the lowest possible clustering error rate that no estimator can uniformly surpass. Notably, $I(p,q)$ is exactly the Renyi divergence (Rényi,, 1961) of order $1/2$ between two Bernoulli distributions, Bern$(p)$ and Bern$(q)$, and hence $\sum_{j\in[J]}I(\theta_{j,k_{1}},\theta_{j,k_{2}})$ can be interpreted as the Renyi divergence between two Bernoulli random vectors with parameters $\mathbf{\Theta}_{:,k_{1}}$ and $\mathbf{\Theta}_{:,k_{2}}$, repsectively. $I^{\star}$ can be regarded as the overall signal-to-noise ratio (SNR) which further takes account into all possible pairs of $(k_{1},k_{2})\in[K]^{2}$ with $k_{1}\neq k_{2}$. Although the form of $I^{\star}$ might seem complicated at its first glance, it actually exactly reflects the likelihood information in the LCMs. By definition of the minimax lower bound, we know that $\exp\left(-I^{\star}/2\right)$ cannot be larger than the spectral clustering error rate $\exp\left(-\Delta^{2}/(8\bar{\sigma}^{2})\right)$ obtained in Proposition 1. We will discuss this in more detail in Section 4.4.
@@ 4.3 Theoretical Guarantee of SOLA
2Initial label estimates: ~$\widetilde{}\mathbf{s}^{(m)}=\textsf{Spec}\left(\mathbf{R}_{{\cal S}_{m}}\right)\in\left\{0,1\right\}^{N/2}$ for $m=1,2$.
@@
$\widehat{s}_{i}^{(m)}=\operatorname*{arg\,max}_{k\in[K]}\sum_{j\in[J]}\Big{[}R_{i,j}\log\Big{(}\widehat{\theta}^{(m)}_{j,k}\Big{)}+(1-R_{i,j})\log\Big{(}1-\widehat{\theta}^{(m)}_{j,k}\Big{)}\Big{]},\quad\forall i\in{\cal S}_{l},\quad l\neq m\in\left\{1,2\right\}.$ (16)
@@
Suppose Assumption 1-2 and the assumptions of Proposition 1 hold. In addition, assume that $N\gtrsim K\log N$, $JK\lesssim\min\left\{\exp\left(-c\Delta^{2}/\bar{\sigma}^{2}\right),N^{\gamma_{0}}\right\}$, for some universal constants $80<c<1/$ 8 and $\gamma_{0}>0$. If the signal-to-noise ratio condition satisfies
@@
Combining Theorem 1 and Theorem 2, we can see that the estimator of Algorithm 3 achieves the fundamental statistical limit and is optimal for estimating the latent class labels. We have several remarks on Theorem 2. First, it is interesting to consider the regime when $J=o\left(\sqrt{\frac{N\log N}{K}}\right)$, since otherwise the signal-to-noise ratio condition would already implies exact recovery (perfect estimation of latent class labels with high probability) for spectral clustering, as shown in Theorem 2. Second, the sample splitting step is crucial in the current proof of Theorem 2 to decouple the dependence between $R_{i,j}$ and ^$\widehat{}\mathbf{\Theta}$ in (8). Alternatively, it may be possible in the future to show that the same theoretical guarantee holds for the original version of SOLA, i.e., Algorithm 2, by developing a leave-one-out analysis similar to that in Zhang and Zhou, (2024). However, The analysis will be highly technical and much more involved than that in Zhang and Zhou, (2024) and beyond the scope of our current focus. Last, although sample splitting simplifies theoretical analysis, it may not be necessary in practice; we use the original SOLA (Algorithm 2) for all numerical experiments and observe satisfactory empirical results.
@@
Our work is related to the Tensor-EM method introduced by Zeng et al., (2023), which also follows an initialization-and-refinement strategy. In the high level, both our method require moment‐based methods in the initialization. However, there are two important distinctions. Zeng et al., (2023) employs a tensor decomposition that leverages higher-order moments and can be computationally intensive and unstable. In contrast, our spectral initialization exploits only the second moment of data (since SVD of the data matrix is equivalent to eigendecomposition of the Gram matrix), making it both simple and efficient. Second, The error rate established in Zeng et al., (2023) is of polynomial order, i.e., $O\bigl{(}J^{-1/2}\bigr{)}$, whereas our SOLA achieves the statistically optimal exponential error rate with respect to $I^{\star}$.
@@ 4.5 Estimation of the Number of Latent Classes $K$
In practice, the true number of latent classes $K$ may be unknown. To see how we can estimate $K$ from data, recall that $\mathbf{R}=\mathbf{Z}\mathbf{\Theta}^{\top}+\mathbf{E}$. By Weyl’s inequality and classical random‐matrix theory (Vershynin,, 2018), the singular values of $\mathbf{R}$ satisfy
@@
This thresholding rule provides a formal analogue of the classical scree‐plot method (Cattell,, 1966; Zhang et al.,, 2020), and its proof shares a similar spirit of Theorem 2 in Zhang et al., (2020) but with a sharper characterization of the constant factor in the noise magnitude.
@@ 5.1 Simulation Studies
Simulation 1: Accuracy for dense $\mathbf{\Theta}$ close to $1/2$.
@@
Figure 2: (Simulation 1) Mis-clustering proportions v.s. number of items $J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(5,5\right)$.
Simulation 2: Accuracy for sparse $\mathbf{\Theta}$
In the second simulation, we choose $(\beta_{1},\beta_{2})=(1,8)$, a sparse scenario where the $\theta_{j,k}$’s tend to be close to 00. Figure 3 illustrates that our proposed methods exhibit superior clustering accuracy when $N$ and $J$ are small, while maintaining competitive performance as these dimensions increase. In this sparse setting, the traditional EM algorithm performs the worst because $\theta_{j,k}$’s are near the boundary (close to 00), making it difficult for random initializations to converge to optimal solutions. Furthermore, the gap between spectral clustering and SOLA becomes more pronounced when the $\theta_{j,k}$’s are bounded away from $0.5$.
Figure 3: (Simulation 2) Mis-clustering proportions v.s. number of items $J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Simulation 3: Stability analysis
@@
Figure 4: (Simulation 3) Failure rate versus the number of items $J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Simulation 4: Computational efficiency
[…]
@@
Table 1: (Simulation 4) Running time (seconds) of different methods.
Figure 5: (Simulation 4) Running time (seconds) of different methods.
5.2 Real-data application

### v2

@@ 4 Theoretical Guarantees
$\theta_{j,k}\in[c_{\theta},C_{\theta}]$ for some universal constants $0<c_{\theta}<C_{\theta}<1$, and $j\in[J]$, $k\in[K]$.
@@ 4.1 Theoretical Guarantee of Spectral Clustering
Proposition 1 (Adapted from Theorem 3.1 in Zhang and Zhou (2024)).
[…]
@@
Proposition 1 delivers that the simple and computationally efficient spectral clustering can essentially lead to exponentially small error rate with respect to the latent class separation $\Delta$. Notably, we can have exact recovery of all latent class labels, (that is, $\ell\left(\widetilde{\bm{s}},{\bm{s}}^{\star}\right)$ with $\widetilde{s}_{i}=\pi(s_{i})$ for all $i=1,\ldots,N$) with high probability as long as $\Delta^{2}/\bar{\sigma}^{2}\geq 8\left(1+c\right)\log N$. In the LCM we study, we have $\bar{\sigma}\leq 1/4$ by definition and hence it suffices to have $\Delta^{2}\geq 2(1+c)\log N$ to achieve exact recovery. This is not hard to achieve even for slowly growing number of items, because if we have constant separation of item parameters for each item with $\delta=\min_{k_{1}\neq k_{2}\in[K],j\in[J]}|\theta_{j,k_{1}}-\theta_{j,k_{2}}|$, then $\Delta^{2}=J\delta^{2}$. In this case, having $J>2(1+c)\delta^{-2}\log N$ for exact recovery corresponds to a moderately growing dimension. In sharp contrast, the JML estimator analyzed in Zeng et al. (2023) only achieves a polynomial convergence rate of order $O\left(J^{-1/2}\right)$. Moreover, they can not achieve exact recovery according to the technical condition and conclusion therein. Therefore, the new spectral estimator $\widetilde{\bm{s}}$ represents a substantial improvement with exponential error decay in (13).
@@
Despite the above nice theoretical guarantee for spectral clustering, the clustering error’s reliance on the Euclidean metric $\Delta$ in Proposition 1 is inherent from the general framework analyzed in Zhang and Zhou (2024), and it turns out to be not statistically optimal due to ignorance of the likelihood information. We will discuss and address this issue next.
@@ 4.2 Fundamental Statistical Limit for LCMs
In this subsection, we provide the understanding of the information-theoretic limit (fundamental statistical limit) of clustering in LCMs. To compare different statistical procedures, we consider the minimax framework that had been widely considered in both statistical decision theory and information theory in the past decades (Le Cam, 2012; Takezawa, 2005). The minimax risk is defined as the infimum, over all possible estimators, of the maximum loss (here, the mis-clustering error) taken over the entire parameter space. In other words, it is the rate of the best estimator one can have among all statistical procedures in the worst-case scenario.
Formally, we establish the minimax lower bound for mis-clustering error under the fixed-effect LCM in the following result, whose proof is adapted from Lyu et al. (2025) and can be found in Section B.1.
@@
Roughly speaking, Theorem 1 informs us that under the fixed-effect LCM, the error rate $\exp\left(-I^{\star}/2\right)$ is the lowest possible clustering error rate that no estimator can uniformly surpass. Notably, $I(p,q)$ is exactly the Renyi divergence (Rényi, 1961) of order $1/2$ between two Bernoulli distributions, Bern(p)(p) and Bern(q)(q), and hence $\sum_{j\in[J]}I(\theta_{j,k_{1}},\theta_{j,k_{2}})$ can be interpreted as the Renyi divergence between two Bernoulli random vectors with parameters $\mathbf{\Theta}_{:,k_{1}}$ and $\mathbf{\Theta}_{:,k_{2}}$, repsectively. $I^{\star}$ can be regarded as the overall signal-to-noise ratio (SNR) which further takes account into all possible pairs of $(k_{1},k_{2})\in[K]^{2}$ with $k_{1}\neq k_{2}$. Although the form of $I^{\star}$ might seem complicated at its first glance, it actually exactly reflects the likelihood information in the LCMs. By definition of the minimax lower bound, we know that $\exp\left(-I^{\star}/2\right)$ cannot be larger than the spectral clustering error rate $\exp\left(-\Delta^{2}/(8\bar{\sigma}^{2})\right)$ obtained in Proposition 1. We will discuss this in more detail in Section 4.4.
@@ 4.3 Theoretical Guarantee of SOLA
2Initial label estimates: $\widetilde{}\mathbf{s}^{(m)}=\textsf{Spec}\left(\mathbf{R}_{{\cal S}_{m}}\right)\in\left\{0,1\right\}^{N/2}$ for $m=1,2$.
@@
$\widehat{s}_{i}^{(m)}=\operatorname*{arg\,max}_{k\in[K]}\sum_{j\in[J]}\Big[R_{i,j}\log\Big(\widehat{\theta}^{(m)}_{j,k}\Big)+(1-R_{i,j})\log\Big(1-\widehat{\theta}^{(m)}_{j,k}\Big)\Big],\qquad\forall i\in{\cal S}_{l},\qquad l\neq m\in\left\{1,2\right\}.$ (16)
@@
Suppose Assumption 1-2 and the assumptions of Proposition 1 hold. In addition, assume that $N\gtrsim K\log N$, $JK\lesssim\min\left\{\exp\left(-c\Delta^{2}/\bar{\sigma}^{2}\right),N^{\gamma_{0}}\right\}$, for some universal constants $0<c<1/8$ and $\gamma_{0}>0$. If the signal-to-noise ratio condition satisfies
@@
Combining Theorem 1 and Theorem 2, we can see that the estimator of Algorithm 3 achieves the fundamental statistical limit and is optimal for estimating the latent class labels. We have several remarks on Theorem 2. First, it is interesting to consider the regime when $J=o\left(\sqrt{\frac{N\log N}{K}}\right)$, since otherwise the signal-to-noise ratio condition would already implies exact recovery (perfect estimation of latent class labels with high probability) for spectral clustering, as shown by Theorem 2 and Proposition 1. Second, the sample splitting step is crucial in the current proof of Theorem 2 to decouple the dependence between $R_{i,j}$ and $\widehat{}\mathbf{\Theta}$ in (8). This step serves purely as a theoretical device to facilitate the theoretical analysis of clustering error rate, as commonly adopted in theoretical statistics (Lei and Zhu, 2017; Vu, 2018; Abbe, 2018; Rinaldo et al., 2019). In practice, halving the sample only increases the estimation error of $\widehat{\bm{\Theta}}$ by a constant factor, which is negligible in theory and immaterial in practice when the sample size $N$ is large. Alternatively, it may be possible in the future to show that the same theoretical guarantee holds for the original version of SOLA, i.e., Algorithm 2, by developing a leave-one-out analysis similar to that in Zhang and Zhou (2024). However, the analysis will be highly technical and much more involved than that in Zhang and Zhou (2024) and beyond the scope of our current focus. Last, although sample splitting simplifies theoretical analysis, it may not be necessary in practice; we use the original SOLA (Algorithm 2) for all numerical experiments and observe satisfactory empirical results.
@@
Our work is related to the Tensor-EM method introduced by Zeng et al. (2023), which also follows an initialization-and-refinement strategy. In the high level, both our method require moment‐based methods in the initialization. However, there are two important distinctions. Zeng et al. (2023) employs a tensor decomposition that leverages higher-order moments and can be computationally intensive and unstable. In contrast, our spectral initialization exploits only the second moment of data (since SVD of the data matrix is equivalent to eigendecomposition of the Gram matrix), making it both simple and efficient. Second, The error rate established in Zeng et al. (2023) is of polynomial order, i.e., $O\bigl(J^{-1/2}\bigr)$, whereas our SOLA achieves the statistically optimal exponential error rate with respect to $I^{\star}$.
@@ 4.5 Estimation of the Number of Latent Classes $K$
In practice, the true number of latent classes $K$ may be unknown. To see how we can estimate $K$ from data, recall that $\mathbf{R}=\mathbf{Z}\mathbf{\Theta}^{\top}+\mathbf{E}$. By Weyl’s inequality and classical random‐matrix theory (Vershynin, 2018), the singular values of $\mathbf{R}$ satisfy
@@
This thresholding rule provides a formal analogue of the classical scree‐plot method (Cattell, 1966; Zhang et al., 2020), and its proof shares a similar spirit of Theorem 2 in Zhang et al. (2020) but with a sharper characterization of the constant factor in the noise magnitude.
@@ 5.1 Simulation Studies
Remark 3 (EM implementations: polca vs. LatentGOLD).
Beyond polca, we also evaluated LatentGOLD (Vermunt and Magidson, 2005), which implements a robust EM routine with safeguards against numerical underflow. Table 1 reports a representative comparison across three regimes ($N=2J$, $N=J$, and $N=J/2$), showing that LatentGOLD is substantially more stable and accurate than polca in these settings. Our overall message remains unchanged: EM-type methods can be competitive in the classical regime with relatively large $N$ and smaller $J$, whereas SOLA/SOLA+ become increasingly advantageous as $J$ grows and the problem enters the high-dimensional regime (with $J$ comparable to or larger than $N$). While retaining polca for comparison purposes in our large-scale simulation studies, we emphasize that when analyzing data with a small-to-moderate $J$, it is recommended to use LatentGOLD for their better numerical stability and performance.
Setting \  Method SOLA SOLA+ EM by polca EM by LatentGold
$N=190,J=95$ ($N=2J$) 0.0815 0.0772 0.3022 0.0645
$N=95,J=95$ ($N=J$) 0.1760 0.1745 0.4642 0.2609
$N=75,J=150$ ($N=J/2$) 0.0411 0.0411 0.5002 0.2795
Table 1: Clustering error across different methods and settings under $100$ replicates
Simulation 1: Latent class estimation accuracy for dense $\mathbf{\Theta}$ close to $1/2$.
@@
Figure 2: Simulation 1: Mis-clustering proportions v.s. number of items $J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(5,5\right)$.
Simulation 2: Latent class estimation accuracy for sparse item parameters $\mathbf{\Theta}$.
In the second simulation, we choose $(\beta_{1},\beta_{2})=(1,8)$, a sparse scenario where the $\theta_{j,k}$’s tend to be close to 0. Figure 3 illustrates that our proposed methods exhibit superior clustering accuracy when $N$ and $J$ are small, while maintaining competitive performance as these dimensions increase. In this sparse setting, the traditional EM algorithm performs the worst because $\theta_{j,k}$’s are near the boundary (close to 0), making it difficult for random initializations to converge to optimal solutions. Furthermore, the gap between spectral clustering and SOLA becomes more pronounced when the $\theta_{j,k}$’s are bounded away from $0.5$.
Figure 3: Simulation 2: Mis-clustering proportions v.s. number of items $J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Simulation 3: Stability analysis.
@@
Figure 4: Simulation 3: Failure rate versus the number of items $J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Simulation 4: Computational efficiency.
Finally, we compare the computation time required by each method. In our experiments, the SVD step and our SOLA and $\text{SOLA}+$ algorithms are implemented in R. The total computation time for Tensor-CEM is the sum of the Matlab initialization and the R refinement stages. As reported in Table 2 and illustrated in Figure 5, our proposed methods are significantly faster than both the EM and Tensor-CEM approaches, with computation time comparable to that of SVD.
SVD SOLA $\text{SOLA}+$ EM Tensor-CEM
@@
Table 2: Simulation 4: Running time (seconds) of different methods.
Figure 5: Simulation 4: Running time (seconds) of different methods.
Simulation 5: Latent class estimation accuracy under different regimes.
In the fifth simulation study, we also employ Spec-EM (EM algorithm initialized by spectral clustering), motivated by the fact that when the number of items is moderately large, spectral clustering provides an effective initialization for the classical EM algorithm for LCMs. In particular, we consider the following regimes: (i) $N=J$ (Figure 6), (ii) $N=J/2$ (Figure 7), (iii) fix $N=100$ and vary $J$ (Figure 8), (iv) fix $J=100$ and vary $N$ (Figure 9), (v) fix $J=30$ and vary $N$ (Figure 10), (vi) $N=J$ with $\left\{s_{i}^{\star}\right\}_{i=1}^{N}$ being independently generated from $[3]$ with probability $(p_{1},p_{2},p_{3})=(1/6,1/3,1/2)$ (Figure 11). In all these settings, we set $(\beta_{1},\beta_{2})=(1,8)$. In addition, we consider (vii) $N=J$ under the fixed design of $\mathbf{\Theta}$ (Figure 12), where the fixed item parameter matrix $\bm{\Theta}$ is constructed by vertically concatenating copies of a $5\times 3$ block matrix $\bm{\Theta}_{0}$ as follows:
$\bm{\Theta}=\begin{bmatrix}\bm{\Theta}_{0}\\ \vdots\\ \bm{\Theta}_{0}\end{bmatrix}\Bigg\}~b~\text{copies},\qquad\bm{\Theta}_{0}:=\begin{bmatrix}0.3&0.7&0.7\\ 0.3&0.7&0.3\\ 0.7&0.3&0.7\\ 0.7&0.3&0.3\\ 0.7&0.7&0.3\end{bmatrix}$.
Here the block matrix $\bm{\Theta}_{0}$ is vertically stacked $b$ times to yield $J=5b$ items with $K=3$ latent classes. It is clear that SOLA and SOLA+ perform significantly better in most scenarios except the most traditional large-$N$, small-$J$ regime, where Spec-EM performs better. This underpins both the validity of SOLA and spectral clustering.
Figure 6: Simulation 5-1: Mis-clustering proportions v.s. number of items $J$ under $N=J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Figure 7: Simulation 5-2: Mis-clustering proportions v.s. number of items $J$ under $N=0.5J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Figure 8: Simulation 5-3: Mis-clustering proportions v.s. number of items $J$ under $N=100$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Figure 9: Simulation 5-4: Mis-clustering proportions v.s. sample size NN under $J=100$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Figure 10: Simulation 5-5: Mis-clustering proportions v.s. sample size NN under $J=30$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Figure 11: Simulation 5-6: Mis-clustering proportions v.s. number of items $J$ under imbalanced latent classes with $N=J$. Entries of $\mathbf{\Theta}$ are independently generated from $\text{Beta}\left(1,8\right)$.
Figure 12: Simulation 5-7: Mis-clustering proportions v.s. number of items $J$ under $N=J$. The $\mathbf{\Theta}$ matrix is set to a fixed design matrix.
Simulation 6: Estimation of number of latent classes.
Our theoretical guarantee for the selection criterion of $K$ relies on random matrix theory under the asymptotic regime $N,J\to\infty$. This means relatively large $N$ and $J$ are needed for the theoretical guarantee to hold, which is consistent with our focus on high-dimensional data. The constant “2” in the threshold reflects a worst-case bound and can indeed be conservative, especially when entries of $\bm{\Theta}$ are mostly concentrated away from $1/2$. To illustrate its effectiveness in practice, we have performed simulation studies with $N=2J$ and $K=3$, under several generative mechanisms for $\bm{\Theta}$ (i.i.d. Beta(1,1)(1,1), Beta(1,4)(1,4), Beta(1,5)(1,5), which reflect different sparsity levels, and also the fixed design as described in Simulation 5). As shown in Table 3, the criterion successfully recovers $K$ in high-dimensional regimes. Our real-data analysis in Section 5.2 also demonstrates that the method works well when the data are sufficiently large. For smaller $N$ and $J$, we recommend that practitioners use classical information criteria, such as BIC, as a more reliable and interpretable tool for selecting $K$.
$J$ $1000$ $1200$ $1400$ $1600$ $1800$ $2000$
Fixed design of item parameters $\% 100100\%$ $\% 100100\%$ $\% 100100\%$
$\theta_{j,k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}$Beta(1,1)(1,1) $\% 100100\%$ $\% 100100\%$ $\% 100100\%$
$\theta_{j,k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}$Beta(1,4)(1,4) $\% 00\%$ $\% 100100\%$ $\% 100100\%$
$\theta_{j,k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}$Beta(1,5)(1,5) $\% 00\%$ $\% 00\%$ $70\%$ $100\%$
Table 3: Proportion of successfully selecting $\widehat{K}$ to be the true $K$ based on $200$ simulation replicates, under different generative mechanisms of the item parameters $\bm{\Theta}=(\theta_{j,k})_{J\times K}$.
5.2 Real Data Example